A 100 pF capacitor is charged to a potential difference of 24 V. It is connected to an uncharged capacitor of capacitance 20 pF. What will be the new potential difference across the 100 pF capacitor?

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. Given that C = 100 PF = 100 * 10^-12 F C base cq = 20 PF = 20 * 10^-12 F V = 24 V q = 24 * 100 * 10^-12 = 24 * 10^-10 Q base 2 = ? Let q base 1 = The new charge 100 PF V base 1 = The Voltage. After the flow of charge, potential is same in the two capacitor V base 1 = q base 2/C base 2 = q base 1/C base 1 = q – q base 1/C base 2 = q base 1/C base 1 = 24 * 10^-10 –q base 1/24 * 10^-12 = q base 1/100 * 10^-12 = 24 * 10^-10 – q base 1 = q base 1/5 = 6q base 1 = 120 * 10^-10 = q base 1 = 120/6 * 10^-10 = 20 * 10^10 ∴ V base 1 = q base 1/ C base 1 = 20 *10^-10/100 * 10^-12 = 20 V